A 25 cm3 sample of an aqueous solution of barium hydroxide, of concentration 0.146 mol dm3 was exactly neutralized by 28.7cm3 of aqueous nitric acid, according to the unbalanced equation:

Ba (OH)2 (aq) + HNO3(aq) ˗˗˗˗˗˗˗˗˃ Ba(NO3)2 (aq) + H2O (l)

i. Calculate the amount of nitric acid in terms molarity and in terms of number of moles


1
Expert's answer
2021-05-13T04:57:17-0400

Ba(OH)2 : 1M=2N

0.146mol dm-3= 0.146M=0.292N

25cm3 of 25ml of 0.292N means 25×0.292= 7.3

HNO3 :1M= 1N

Conc of HNO3 = x(N)

28.7 mol of x(N) means 28.7×x mili equFor exact neutralization 28.7×x= 7.3

X= 0.25

Concentration of nitric acid is 0.254M= 0.254 mol dm-3